OpenStudy (anonymous):

Find the distance between each pair of points: a(4,7) b(1,3)

OpenStudy (anonymous):

I think you would do sqrt((x1-x2)^2+(y1+y2)^2)=distance. The ones and twos next to the x and y are supposed to be subscripts.

OpenStudy (anonymous):

oh sorry sqrt((x1-x2)^2+(y1-y2)^2)=distance I accidentally wrote a plus btw the y's

OpenStudy (anonymous):

\[\sqrt{(1-4)^2+(3-7)^2}\] \[\sqrt{(-3)^2+(-4)^2}\] \[\sqrt{9+16}\] \[\sqrt{25}\] 5